# A systematic search for big dippers in ASAS-SN

### v1

@@ 4. Discussion
Here, we only searched for long-period binaries and dipper stars within the magnitude range of mag. Additional searches across the entire ASAS-SN magnitude range ( mag) would certainly detect many more sources. This search is also complementary to the search for slowly varying ASAS-SN stars in Petz & Kochanek (2025) and Petz et al. (2025). There, seasonal medians were used to eliminate shorter time scale variability and then linear and quadratic fits to the medians were used to identify sources slowly varying in brightness. Of the sources here, J080327-261620 (ASASSN-24fa) was also flagged in Petz & Kochanek (2025).
@@
| Source | RA | DEC | Search Method | Number of Dips | Maximum Duration | Start of Most Recent Dip | Maximum Depth |
| [deg] | [deg] | [days] | [mag] | ||||
| Single Eclipse Binaries | |||||||
| J073234200049 | Known | 07/10/2023 | |||||
| J175602013135 | Known | 22/02/2024 | |||||
| J223332565552 | Known | 15/11/2023 | |||||
| J073924272916 | Pipeline | 25/02/2023 | |||||
| J074007161608 | Pipeline | 21/09/2022 | |||||
| J094848545959 | Pipeline | 20/03/2019 | |||||
| J162209444247 | Pipeline | 25/04/2023 | |||||
| J183210173432 | Pipeline | 02/09/2018 | |||||
| J183606314826 | Pipeline | 10/03/2022 | |||||
| J190316195739 | Pipeline | 19/08/2019 | |||||
| J205245713514 | Pipeline | 14/04/2023 | |||||
| J212132480140 | Pipeline | 09/09/2021 | |||||
| J225702562312 | Pipeline | 08/06/2019 | |||||
| Multiple Eclipse Binaries | |||||||
| J005437644347 | Known | 01/12/2024 | |||||
| J181752580749 | Known | 07/08/2024 | |||||
| J005355633515 | Pipeline | 06/08/2020 | |||||
| J062510075341 | Pipeline | 14/08/2019 | |||||
| J124745622756 | Pipeline | 14/05/2024 | |||||
| J160757574540 | Pipeline | 17/04/2024 | |||||
| J175912120956 | Pipeline | 12/05/2022 | |||||
| Dippers | |||||||
| J070519061219 | Known | 23/08/2024 | |||||
| J081523385923 | Known | 30/03/2021 | |||||
| J085816430955 | Known | Ongoing | 06/04/2025 | ||||
| J114712621057 | Known | 20/11/2020 | |||||
| J181707164819 | Known | Ongoing | 11/05/2025 | ||||
| J183153284827 | Known | Ongoing | 06/03/2025 | ||||
| J184916473251 | Known | 08/07/2021 | |||||
| J042214152530 | Pipeline | 04/10/2022 | |||||
| J080327261620 | Pipeline | Ongoing | 31/01/2022 | ||||
| J174328343315 | Pipeline | 25/02/2020 | |||||
| J202402383938 | Pipeline | 15/04/2018 | |||||
@@ Acknowledgements
DMR acknowledges support from the OSU Presidential Fellowship. CSK, KZS, and DMR are supported by NSF grants AST-2307385 and 2407206. B.J.S is supported with funds from NSF (grant AST-2407205) and NASA (grants HST-GO-17087, 80NSSC24K0521, 80NSSC24K0490, 80NSSC23K1431). S.D. acknowledges the National Natural Science Foundation of China (Grant No. 12133005) and the New Cornerstone Science Foundation through the XPLORER PRIZE.

### v2

@@ 4. Discussion
Here, we only searched for long-period binaries and dipper stars within the magnitude range of mag. Additional searches across the entire ASAS-SN magnitude range ( mag) would certainly detect many more sources. This search is also complementary to the search for slowly varying ASAS-SN stars in Petz and Kochanek (2025) and Petz et al. (2025). There, seasonal medians were used to eliminate shorter time scale variability and then linear and quadratic fits to the medians were used to identify sources slowly varying in brightness. Of the sources here, J080327-261620 (ASASSN-24fa) was also flagged in Petz and Kochanek (2025).
@@
| Source | RA | DEC | Search Method | Number of Dips | Maximum Dip Duration | Minimum Period | Start of Most Recent Dip | Maximum Depth |
| [deg] | [deg] | [days] | [days] | [mag] | ||||
| Single Eclipse Binaries | ||||||||
| J073234200049 | Known | 40.9 | 07/10/2023 | |||||
| J175602013135 | Known | 94.8 | 22/02/2024 | |||||
| J223332565552 | Known | 19.9 | 15/11/2023 | |||||
| J073924272916 | Pipeline | 26.9 | 25/02/2023 | |||||
| J074007161608 | Pipeline | 4.0 | 21/09/2022 | |||||
| J094848545959 | Pipeline | 7.7 | 20/03/2019 | |||||
| J162209444247 | Pipeline | 3.1 | 25/04/2023 | |||||
| J183210173432 | Pipeline | 41.0 | 02/09/2018 | |||||
| J183606314826 | Pipeline | 10.7 | 10/03/2022 | |||||
| J190316195739 | Pipeline | 15.3 | 19/08/2019 | |||||
| J205245713514 | Pipeline | 2.0 | 14/04/2023 | |||||
| J212132480140 | Pipeline | 57.8 | 09/09/2021 | |||||
| J225702562312 | Pipeline | 6.0 | 08/06/2019 | |||||
| Multiple Eclipse Binaries | ||||||||
| J005437644347 | Known | 10.9 | 01/12/2024 | |||||
| J181752580749 | Known | 86.8 | 07/08/2024 | |||||
| J005355633515 | Pipeline | 10.1 | 06/08/2020 | |||||
| J062510075341 | Pipeline | 4.3 | 14/08/2019 | |||||
| J124745622756 | Pipeline | 7.2 | 14/05/2024 | |||||
| J160757574540 | Pipeline | 8.3 | 17/04/2024 | |||||
| J175912120956 | Pipeline | 0.0 | 12/05/2022 | |||||
| Dippers | ||||||||
| J070519061219 | Known | 274.8 | 23/08/2024 | |||||
| J081523385923 | Known | 608.1 | 30/03/2021 | |||||
| J085816430955 | Known | ¿ 200 days, Ongoing | 06/04/2025 | |||||
| J114712621057 | Known | 1282.8 | 20/11/2020 | |||||
| J181707164819 | Known | ¿ 200 days, Ongoing | 11/05/2025 | |||||
| J183153284827 | Known | ¿ 200 days, Ongoing | 06/03/2025 | |||||
| J184916473251 | Known | 273.1 | 08/07/2021 | |||||
| J042214152530 | Pipeline | 44.2 | 04/10/2022 | |||||
| J080327261620 | Pipeline | ¿ 1400 days, Ongoing | 31/01/2022 | |||||
| J174328343315 | Pipeline | 22.1 | 25/02/2020 | |||||
| J202402383938 | Pipeline | 1709.6 | 15/04/2018 | |||||
@@ Acknowledgements
We thank the anonymous reviewer for their comments. DMR acknowledges support from the OSU Presidential Fellowship. CSK, KZS, and DMR are supported by NSF grants AST-2307385 and 2407206. B.J.S is supported with funds from the NSF (grant AST-2407205). S.D. acknowledges the National Natural Science Foundation of China (Grant No. 12133005) and the New Cornerstone Science Foundation through the XPLORER PRIZE.